OpenStudy (anonymous):

simplify: f(x)=9-xpower2 f(8)= f(-9) f(k)= f(kpower2-7)=

OpenStudy (anonymous):

well f(x)=9-x^2 is that written better. then f(8)=9-8^2=9-64=-55

OpenStudy (anonymous):

f(-9)=9-(-9)^2=9-81=-72

OpenStudy (anonymous):

I'm a little confused about the f(k) part. could you rewrite it?

OpenStudy (anonymous):

f((k^2)- 7 ) = 9 - ((k^2)- 7 )^2 = 9 - ( k^4 -14k^2 +49) = 9 - k^4 + 14k^2 - 49

OpenStudy (anonymous):

gotcha. haha wasn't thinking about that. good teamwork.

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

f(k)=